Is this solvable using the Lagrange method? If so, Does anyone know how to solve this using the Lagrange method? Utility is alnx1 + lnx2. Budget constraint is P1x1+P2x2= W


1
Expert's answer
2020-11-09T10:22:42-0500

Utility function given as below

"U(x_1, x_2)= a \ln x_1 + \ln x_2"

taking the differentiation of utility function,

"dU=\dfrac{adx_1}{x_1}+\dfrac{dx_2}{x_2}=0"

"\Rightarrow \dfrac{dx_1}{dx_2}=\frac{-x_1}{ax_2}"

Budget constraint is given as

"P_1x_1+P_2x_2= W"

We will substitute the values in this we can proceed for the solution.

As it is not a complete question so the mentioned instruction is sufficient to solve this type question.